Find a polynomial, P(x), of degree 3 with zeros of 4,1 and −1, if P(0) = 8

1
Expert's answer
2022-04-06T16:12:50-0400

"P(x)=a(x-4)(x-1)(x+1)".

"P(0)=8 \to a(-4)(-1)*1=8 \to a=2."

"P(x)=2(x-4)(x-1)(x+1)=2x^3-8x^2-2x+8."